OpenStudy (anonymous):

Evaluate the expression 3(7+4) squared 2-14÷7

alones (alones):

\( \color{eneter color name}{3(7+4) ^2-14÷7 } \)

OpenStudy (vhtran):

Use PEMDAS, so first what is 7+4?

OpenStudy (vhtran):

\[3 \times 11^2-14\div7\]

OpenStudy (vhtran):

\[11^2=121\]

OpenStudy (vhtran):

\[3 \times 121 =363\]

OpenStudy (vhtran):

14/7 = 2

OpenStudy (vhtran):

363-2 = 361